OpenStudy (anonymous):

Solve 3x2 = -30x + 24. solve me though this please?

OpenStudy (anonymous):

answer choices: quantity of negative 10 plus or minus square root of 33 all over 2 negative 5 plus or minus 2 square root of 33 negative 10 plus or minus square root of 33 negative 5 plus or minus square root of 33

OpenStudy (uri):

http://blogs.discovermagazine.com/loom/files/2008/07/quadratic.jpg Quadratic eq and a:3 b:30 c:-24

OpenStudy (uri):

3x2 = -30x + 24. 3x^2+30x-24=0

OpenStudy (anonymous):

\[3x^2=-30+24\]Divide everything by 3.\[x^2=-10+8\]Put everything on one side.\[x^2+10-8=0\] Let's use the quadratic formula.\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]\[\frac{ -10 \pm \sqrt{10^2-4(1)(-8)} }{ 2(1) }\]\[\frac{ -10 \pm \sqrt{132} }{ 2 }\]\[\frac{ -10 \pm 2\sqrt{33} }{ 2 }\] \[-5 \pm \sqrt{33}\] The last answer choice is correct.

OpenStudy (anonymous):

okay thanks!